please answer this question:
a car is travelling at the rate of 60 km. per hour while another car is travelling at the rate of 7 km per hour.in how many hours will the two cars be when they are 260km. apart?

Answer by mananth(16946)


I have assumed Car B rate as 70 km/hour instead of 7 which looks practical. If not then replace 70 with 7 and you will get the desired answer.
Car A 60 km/hour
car B 70 km/hour
..
Let Car A travel x km in the time car b travels 260-x
x/60 = 260-x/70
70x=60(260-x)
70x=15600-60x
70x+60x=15600
130x= 15600
x= 120 km
speed of car A = 60 km/hour
120/60 = time
=2 hours.
They will be 260 km apart after 2 hours